# FARM: Functional Group-Aware Representations for Small Molecules

### v3

@@ 5 Experimental Results
Baselines   We consider works that incorporate functional group information to enhance representation learning (Zhang et al., 2020, 2021; Yang et al., 2022; Wang et al., 2023; Han et al., 2023; Li et al., 2023; Zang et al., 2023; Wang et al., 2022b; Xia et al., 2023; Nguyen et al., 2024b), and other approaches that utilize masked atom prediction as a self-supervised training task (Rong et al., 2020; Hu et al., 2019; Liu et al., 2021; Fang et al., 2022; Zhou et al., 2023).
Results   Table 2 and 3 present the performance of FARM alongside other baseline models on seven MoleculeNet classification tasks. FARM consistently outperforms baseline models on various benchmarks, demonstrating its robustness and versatility. Its strong performance across various classification and regression tasks indicates that FARM is a highly effective pretrained model, well-suited for a broad range of downstream tasks in molecular property prediction.
Ablation Study on FARM Component Analysis   To assess the effectiveness of each component in our architecture, we conducted a comprehensive ablation study across several MoleculeNet benchmark tasks. The results are shown in Table 10. For more details about the baselines, please refer to appendix F. The results show that each component of FARM contributes to improved performance compared to the baseline. Combining all components leads to strong overall performance of FARM across all test tasks.
@@
Ablation Study on Fragmentation Methods   We compare our FG-based fragmentation method with BRICS fragmentation. When treating each BRICS fragment as a functional group, the enriched SMILES vocabulary size reaches 575,612 (min freq $\leq$ 5), compared to just 14,741 with our FG-based method. This is because BRICS produces larger, less reusable fragments, resulting in a bloated vocabulary with many rare tokens—making training harder and increasing the risk of out-of-vocabulary issues. As shown in Table 4, our FG-based method consistently outperforms BRICS across all six MoleculeNet tasks under the same setup.
@@ 6 Conclusion, Limitations, and Future Work
In summary, FARM demonstrates robust performance across various MoleculeNet classification tasks, outperforming or matching baseline models. The integration of functional group information and the alignment of FG-enhanced SMILES representations with FG graph embeddings through contrastive learning significantly enhance its effectiveness. While FARM shows strong performance, there is a key limitation that should be addressed in future work. The current model does not incorporate a full 3D molecular representation, which is critical for capturing stereochemistry and spatial configurations that influence molecular properties. Integrating 3D structural information Yan et al. (2024) could further improve the model’s ability to make accurate predictions.
@@
Acknowledgements
This work is supported by the Molecule Maker Lab Institute: an AI research institute program supported by NSF under award No. 2019897. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied of, the National Science Foundation, and the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for governmental purposes notwithstanding any copyright annotation therein.
@@ D.1 Training Masked Language Model for SMILES Representation
Additional details of the training setup include training the BERT model on 20 million SMILES for 15 epochs using two NVIDIA Tesla V100 GPUs. The learning rate was set to $1e-5$ , with a batch size of 128, and model checkpoints were saved after every 10,000 batches. This setup was also applied to the baseline model, which used conventional SMILES for comparison.
@@ D.4 Contrastive Learning: Align SMILES and Structure Representation
[…]
@@
  • $N$ is the number of training examples (or contrastive pairs)
@@ D.5 Downstream task finetuning
MoleculeNet tasks are treated as downstream tasks for our FARM model. We freeze all layers of FARM and pair it with a GRU head for both classification and regression tasks. For classification, we use cross-entropy as the loss function, while for regression, we employ mean squared error. The Adam optimizer is applied with a learning rate of $1e-4$ and a cosine annealing learning rate schedule with a period of 20 epochs. The training process spans 100 epochs with a batch size of 16, using an 80-10-10 train-validation-test split with scaffold splitting. To address imbalanced datasets, we implement a weighted loss function, assigning a weight of 5 to classes with fewer samples. For each task, we conduct three runs with different train-validation-test splits and report the average and standard deviation of the results.
Appendix E Evaluation of FARM on Quantum Mechanical Property Prediction
Table 9: Performance of FARM on quantum mechanical property prediction using the Quantum Mechanical dataset. $R^{2}$ scores are reported with standard deviation over 10 runs.
Class $\mathbf{R^{2}}$ Class $\mathbf{R^{2}}$
HOMOm1 (eV) 0.927±0.01480.927$\pm$ 0.01480.927 ± 0.0148 T5 0.905±0.03150.905$\pm$ 0.03150.905 ± 0.0315
HOMO (eV) 0.895±0.02980.895$\pm$ 0.02980.895 ± 0.0298 S1 0.936±0.01610.936$\pm$ 0.01610.936 ± 0.0161
LUMO (eV) 0.928±0.009480.928$\pm$ 0.009480.928 ± 0.00948 S2 0.948±0.01560.948$\pm$ 0.01560.948 ± 0.0156
LUMOp1 (eV) 0.922±0.02800.922$\pm$ 0.02800.922 ± 0.0280 S3 0.911±0.008490.911$\pm$ 0.008490.911 ± 0.00849
DipoleMoment (Debye) 0.755±0.05320.755$\pm$ 0.05320.755 ± 0.0532 S4 0.919±0.007920.919$\pm$ 0.007920.919 ± 0.00792
LogP 0.995±0.005660.995$\pm$ 0.005660.995 ± 0.00566 S5 0.924±0.01430.924$\pm$ 0.01430.924 ± 0.0143
PrimeExcite (eV) 0.877±0.04910.877$\pm$ 0.04910.877 ± 0.0491 O1 0.839±0.03400.839$\pm$ 0.03400.839 ± 0.0340
PrimeExcite (osc) 0.837±0.02430.837$\pm$ 0.02430.837 ± 0.0243 TDOS2.0 0.868±0.09900.868$\pm$ 0.09900.868 ± 0.0990
T1 0.929±0.005870.929$\pm$ 0.005870.929 ± 0.00587 TDOS2.5 0.877±0.004100.877$\pm$ 0.004100.877 ± 0.00410
T2 0.925±0.01820.925$\pm$ 0.01820.925 ± 0.0182 TDOS3.0 0.815±0.04470.815$\pm$ 0.04470.815 ± 0.0447
T3 0.900±0.01010.900$\pm$ $0$ 0.0101 TDOS3.5 0.895±0.03270.895$\pm$ 0.03270.895 ± 0.0327
T4 0.921±0.01960.921$\pm$ 0.01960.921 ± 0.0196 TDOS4.0 0.919±0.03130.919$\pm$ 0.03130.919 ± 0.0313
TDOS4.5 0.938±0.01830.938$\pm$ 0.01830.938 ± 0.0183
Table 9 presents the results of FARM on the quantum mechanical property prediction tasks from the dataset introduced in [Angello et al., 2024]. FARM demonstrates consistently high performance across a wide range of properties, including difficult-to-predict targets such as Dipole Moment, excited-state energies (S1–S5 and T1–T5), and total density of states (TDOS). These results underscore FARM’s ability to accurately model complex quantum mechanical behaviors by leveraging functional group information. Its success in capturing nuanced molecular interactions further illustrates FARM’s applicability to real-world challenges in chemical property prediction, where understanding the role of functional modules is essential for robust and generalizable predictions.
Appendix F Ablation Study
F.1 FARM Component Analysis
To assess the effectiveness of each component in our architecture, we conducted a comprehensive ablation study across several MoleculeNet benchmark tasks. The first model, FM_KGE + GAT, utilizes FG knowledge graph embeddings as input for a Graph Attention Network [Veličković et al., 2017] (GAT) to predict molecular properties. Although its performance on these tasks is not the strongest, the model still demonstrates its capacity to learn underlying chemical rules (syntax and semantics) from the data to a certain degree.
@@ F.1 FARM Component Analysis
BBBP BACE HIV Average ESOL FreeSolv Average QM9
#tasks 1 1 1 1 1 3
#samples 2039 1513 41127 1128 642 133885
Metric ROC-AUC ($\uparrow$) RMSE ($\downarrow$) MAE ($\downarrow$)
FG_KGE + GAT 73.23 $\pm$ 1.93 76.44 $\pm$ 1.27 71.65 $\pm$ 0.98 73.77 2.35 $\pm$ 0.210 4.32 $\pm$ 0.29 3.335 0.0139 $\pm$ 0.00014
AttentiveFP 77.71 $\pm$ 1.30 77.15 $\pm$ 0.78 78.81 $\pm$ 0.99 77.89 1.63 $\pm$ 0.042 2.11 $\pm$ 0.94 1.87 0.0056 $\pm$ 0.00012
FG AttentiveFP 85.57 $\pm$ 1.32 87.30 $\pm$ 0.90 81.21 $\pm$ 0.92 84.5 1.02 $\pm$ 0.034 1.08 $\pm$ 0.14 1.05 0.0053 $\pm$ 0.00034
BERT 82.12 $\pm$ 1.45 85.12 $\pm$ 0.76 83.03 $\pm$ 1.12 83.42 1.45 $\pm$ 0.056 1.89 $\pm$ 0.09 1.67 0.0059 $\pm$ 0.00012
FG BERT 94.36 $\pm$ 0.50 94.54 $\pm$ 0.40 81.93 $\pm$ 1.70 90.27 0.608 $\pm$ 0.031 0.507 $\pm$ 0.03 0.558 0.0041 $\pm$ 0.00017
FARM 96.23 $\pm$ 0.7 96.19 $\pm$ 0.65 82.13 $\pm$ 1.10 91.43 0.734 $\pm$ 0.039 0.308 $\pm$ 0.08 0.521 0.0038 $\pm$ 0.00014
F.2 Fragmentation Methods

### v4

@@ 5.2 Baselines
MoleculeNet We consider works that incorporate functional group information to enhance representation learning (Zhang et al., 2020; 2021; Yang et al., 2022; Wang et al., 2023; Han et al., 2023; Li et al., 2023; Zang et al., 2023; Wang et al., 2022b; Xia et al., 2023; Nguyen et al., 2024b; Kengkanna and Ohue, 2024), and other approaches that utilize masked atom prediction as a self-supervised training task (Rong et al., 2020; Hu et al., 2019; Liu et al., 2021; Fang et al., 2022; Zhou et al., 2023; Soares et al., 2024).
D-B-A Oligomer Photostability For this dataset, we do not find any prior benchmark studies, as most foundation models for small molecules have focused exclusively on MoleculeNet. We include this dataset to evaluate our model on a diverse set of tasks and to assess its generalizability beyond standard benchmarks.
5.3 Results
MoleculeNet Table 2 presents the performance of FARM alongside other baseline models on 13 MoleculeNet tasks. FARM consistently outperforms baseline models across the benchmarks, demonstrating its robustness and versatility. Notably, FARM achieves particularly strong performance on BBBP and BACE, which are often more learnable because these tasks involve well-defined functional group patterns associated with blood–brain barrier penetration and beta-secretase inhibition, respectively—patterns that FG-aware tokenization captures effectively. FARM also excels on solubility-related tasks such as Lipo, FreeSolv, and ESOL, likely due to its ability to model atom-level chemical context and functional groups that strongly influence solubility. Overall, its strong performance across both classification and regression tasks highlights FARM as a highly effective pretrained model, well-suited for a broad range of downstream applications in molecular property prediction.
D-B-A Oligomer Photostability We evaluate FARM on the D–B–A oligomer photostability dataset across 25 property prediction tasks (results shown in Table 3). The model achieves a lowest R2 of 0.815 for TDOS 3.0 and a highest R2 of 0.995 for LogP, with 16 out of 25 tasks exceeding R2 values of 0.9. These tasks are particularly challenging not only due to the hard-case scaffold split, in which the training set contains no molecules sharing donor or acceptor blocks with the validation and test sets, but also due to the nature of the targets, including difficult-to-predict properties such as dipole moment, excited-state energies (S1–S5 and T1–T5), and total density of states (TDOS). Despite these challenges, FARM demonstrates consistently high performance across all properties. These results highlight FARM’s ability to accurately model complex quantum mechanical behaviors by leveraging functional group information and capturing nuanced molecular interactions, underscoring its robustness and generalizability for real-world chemical property prediction tasks.
5.4 Ablation Study on FARM Component Analysis
To assess the contribution of each component in our architecture: FG-enhanced SMILES modeling via BERT, FG embeddings from the FG knowledge graph, and molecule backbone modeling using FG graphs with GNN link prediction, we conducted a comprehensive ablation study across several MoleculeNet benchmark tasks (Table 4). The results demonstrate that each component of FARM improves performance over the baseline, and the combination of all components yields the strongest overall performance across all test tasks. Specifically, incorporating FG-aware tokenization (FG BERT vs. BERT) leads to a substantial improvement across all tasks, highlighting the importance of injecting chemical context into the molecular representation. Further integrating molecular structure via contrastive learning (FARM vs. FG BERT) consistently achieves the highest performance across 6 out of 7 downstream tasks, demonstrating the complementary nature of the two representations. Note that random splitting is used consistently across models in this ablation study, while scaffold splitting is applied for benchmarking in Tables 2 to ensure fair comparison with other methods.
Model Description BBBP BACE HIV Avg ESOL FreeSolv Avg QM9
#tasks 1 1 1 1 1 3
#samples 2039 1513 41127 1128 642 133885
Metric ROC-AUC ($\uparrow$) RMSE ($\downarrow$) MAE ($\downarrow$)
FG_KGE + GAT FG knowledge graph embeddings + GAT 73.23±1.93 76.44±1.27 71.65±0.98 73.77 2.35±0.210 4.32±0.29 3.335 0.0139±0.00014
AttentiveFP GNN with masked atom prediction (atom type) 77.71±1.30 77.15±0.78 78.81±0.99 77.89 1.63±0.042 2.11±0.94 1.87 0.0056±0.00012
FG AttentiveFP GNN with masked atom prediction (atom type + FG) 85.57±1.32 87.30±0.90 81.21±0.92 84.5 1.02±0.034 1.08±0.14 1.05 0.0053±0.00034
BERT BERT on canonical SMILES 82.12±1.45 85.12±0.76 83.03±1.12 83.42 1.45±0.056 1.89±0.09 1.67 0.0059±0.00012
FG BERT BERT on FG-enhanced SMILES 94.36±0.50 94.54±0.40 81.93±1.70 90.27 0.608±0.031 0.507±0.03 0.558 0.0041±0.00017
FARM FG BERT + GNN link prediction + contrastive learning 96.23±0.7 96.19±0.65 82.13±1.10 91.43 0.734±0.039 0.308±0.08 0.521 0.0038±0.00014
5.5 Ablation Study on Fragmentation Methods
@@
We compare our FG-based fragmentation method with BRICS fragmentation. When treating each BRICS fragment as a functional group, the FG-enhanced SMILES vocabulary expands to 575,612 tokens (min frequency $\leq$ 5), compared to just 14,741 tokens with our chemically grounded FG detection method. This difference arises because BRICS generates larger, less reusable fragments, leading to a bloated vocabulary with many rare tokens, which complicates training and increases the risk of out-of-vocabulary issues. As shown in Table 5, our FG-based method consistently outperforms BRICS across all six MoleculeNet tasks under the same setup.
@@
5.6 Benchmarking on ADMET Tasks
To further demonstrate FARM’s generalization beyond MoleculeNet, we evaluate it on a subset of reliable ADMET tasks from the Therapeutics Data Commons (TDC) leaderboard [^1] https://tdcommons.ai/benchmark/admet_group/overview/, covering Absorption, Distribution, Metabolism, Excretion, and Toxicity properties. We adopt the standardized train/validation/test split from the leaderboard for consistency. As shown in Table 6, FARM achieves state-of-the-art results on 3 out of 7 tasks and demonstrates competitive performance on the remaining tasks. Notably, none of the foundation models included in Table 2 have reported results on these ADMET tasks. We note that many TDC ADMET tasks are limited by small and highly imbalanced datasets, often containing fewer than 100 positive examples, making it difficult to learn generalizable patterns. We therefore evaluate on a subset of more reliable tasks, comparing against all leading models including those whose code is not publicly available. Many leaderboard approaches rely on hand-crafted features and complex pipelines, which are orthogonal to our focus on end-to-end transferable representation learning; combining FARM with such strategies is a promising direction for future work.
Table 6: FARM’s performance on ADMET tasks from the TDC leaderboard. Bold indicates state-of-the-art.
  Dataset   Task   Unit   Metric   Type   SOTA   FARM   Rank
  Bioav   Absorption   %   AUROC   Binary   0.753   0.709   5
  AqSol   log mol/L   MAE   Regression   0.761   0.739   1
  PPBR   Distribution   %   MAE   Regression   7.526   7.376   1
  VDss   L/kg   Spearman   Regression   0.724   0.652   4
  CYP2C9 Inhibition   Metabolism   %   AUPRC   Binary   0.859   0.798   4
  CYP3A4 Inhibition   %   AUPRC   Binary   0.916   0.877   5
  Ames   Excretion   %   AUROC   Binary   0.871   0.875   1
@@ 6 Conclusion, Limitations, and Future Work
In summary, FARM demonstrates robust performance across various MoleculeNet classification tasks, outperforming or matching baseline models. The integration of functional group information and the alignment of FG-enhanced SMILES representations with FG graph embeddings through contrastive learning significantly enhance its effectiveness. While FARM shows strong performance, there is a key limitation that should be addressed in future work. The current model does not incorporate a full 3D molecular representation, which is critical for capturing stereochemistry and spatial configurations that influence molecular properties. Integrating 3D structural information (Yan et al., 2024) could further improve the model’s ability to make accurate predictions.
@@ A.1 Training data
To ensure data quality, we applied standard preprocessing steps: molecules that failed RDKit processing (i.e., invalid SMILES) were removed, duplicate molecules were identified and removed by converting all SMILES to their canonical form using RDKit, and any overlap with the MoleculeNet benchmark datasets used for downstream evaluation was excluded to prevent data leakage.
@@ D.1 Training Masked Language Model for SMILES Representation
Additional details of the training setup include training the BERT model on 20 million SMILES for 15 epochs using two NVIDIA Tesla V100 GPUs. The learning rate was set to 1e51e-5 , with a batch size of 128, and model checkpoints were saved after every 10,000 batches. This setup was also applied to the baseline model, which used conventional SMILES for comparison.
@@ D.4 Contrastive Learning: Align SMILES and Structure Representation
$\text{cosine\_similarity}(\mathbf{u},\mathbf{v})=\frac{\mathbf{u}\cdot\mathbf{v}}{\|\mathbf{u}\|\|\mathbf{v}\|}$
@@
  • NN is the number of training examples (or contrastive pairs)
@@
The overall model architecture consists of three main modules that require training: a BERT model, a FG knowledge graph, and GNN. The FG knowledge graph is represented using a ComplEx model and trained first to generate FG embeddings. These embeddings are then used to train the GNN on a link prediction task. Once the GNN training converges, molecular structure embeddings are extracted and a small projection layer is trained to align them with BERT’s representations via contrastive learning. This alignment is optimized jointly with the masked language modeling loss and the contrastive loss, updating both the BERT model and the projection layer.
@@ D.5 Downstream task finetuning
MoleculeNet tasks are treated as downstream tasks for our FARM model. We freeze all layers of FARM and pair it with a GRU head for both classification and regression tasks. For classification, we use cross-entropy as the loss function, while for regression, we employ mean squared error. The Adam optimizer is applied with a learning rate of 1e41e-4 and a cosine annealing learning rate schedule with a period of 20 epochs. The training process spans 100 epochs with a batch size of 16, using an 80-10-10 train-validation-test split with scaffold splitting. To address imbalanced datasets, we implement a weighted loss function, assigning a weight of 5 to classes with fewer samples. For each task, we conduct three runs with different train-validation-test splits and report the average and standard deviation of the results.
Appendix E Ablation Study
E.1 FARM Component Analysis
To assess the effectiveness of each component in our architecture, we conducted a comprehensive ablation study across several MoleculeNet benchmark tasks. The first model, FG_KGE + GAT, utilizes FG knowledge graph embeddings as input for a Graph Attention Network [Veličković et al., 2017] (GAT) to predict molecular properties. Although its performance on these tasks is not the strongest, the model still demonstrates its capacity to learn underlying chemical rules (syntax and semantics) from the data to a certain degree.
@@ E.1 FARM Component Analysis
E.2 Fragmentation Methods
@@
Appendix F Benchmarking on ADMET Tasks
Table 6 presents the performance of FARM on the ADMET datasets. The ADMET leaderboard [^2] https://tdcommons.ai/benchmark/admet_group/overview/ evaluates model performance across ADMET tasks (Absorption, Distribution, Metabolism, Excretion, Toxicity) using a standardized train/validation/test split, which we adopt for consistency. FARM achieves state-of-the-art results on 3 out of 7 tasks and demonstrates competitive performance with other top models on the remaining tasks, highlighting its robustness in ADMET prediction. Notably, none of the foundation models for small molecules included in Table 2 have reported results on these ADMET tasks, primarily due to the limitations of the datasets themselves.
Many ADMET tasks in TDC are limited by small and highly imbalanced datasets, often containing fewer than 100 positive examples. These constraints make it difficult for models to learn meaningful, generalizable patterns, particularly for complex endpoints like toxicity or clearance, which involve diverse and mechanistically unrelated biological pathways. For instance, a dataset labeling 1,000 compounds as “toxic” may include compounds that exert toxicity through entirely different mechanisms, preventing models from reliably capturing features that generalize to unseen cases. Consequently, performance on many TDC ADMET tasks is often near random, limiting their utility for assessing generalization. Specific tasks such as HIA_Hou, Clearance_Hepatocyte_AZ, and CYP3A4_Substrate_CarbonMangels exemplify this limitation, where even top-performing models fail to achieve meaningful results. To mitigate this, we evaluated FARM on a subset of more reliable TDC ADMET tasks, comparing its performance against all leading models, including some whose code is not publicly runnable (e.g., MapLight models). Many leaderboard approaches rely heavily on hand-crafted features and complex pipelines, which are orthogonal to our focus on end-to-end, transferable representation learning; combining FARM with such strategies presents a promising direction for future work.